Consider the resonance structures for the carbonat
Answer #1
Concepts and reason

Carbonate ion acts as a base and it can form three resonance structures. If a compound has more number of resonance structures then it is more stable.

Fundamentals

Negative charge on each oxygen atom can be calculated by dividing the charge present oxygen atoms with number of oxygen atoms.

Calculate the bond order of each C-O bond by using the following expression.

Bondorder=TotalnumberofbondsNumberofCObonds{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total}}\,{\rm{number}}\,{\rm{of}}\,{\rm{bonds}}}}{{{\rm{Number}}\,{\rm{of}}\,{\rm{C - O}}\,{\rm{bonds}}}}

Part a

The resonance structures of carbonate ion are,

O:
:O
O:
O:
O
I..
:O:
I..
O
O:

Here, two oxygen atoms have a charge of 1 - 1 and other one is neutral.

So,thenegativechargeoneachoxygenatom=2charge3=0.67\begin{array}{c}\\{\rm{So,}}\,{\rm{the}}\,{\rm{negative}}\,{\rm{charge}}\,{\rm{on}}\,{\rm{each}}\,{\rm{oxygen}}\,{\rm{atom}}\,{\rm{ = }}\,\frac{{ - {\rm{2}}\,{\rm{charge}}}}{{\rm{3}}}\\\\ & & & & & \,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\, - {\rm{0}}{\rm{.67}}\\\end{array}

Part b

The resonance structures of carbonate ion are,

Total number of bonds present in the carbonate ion is four.

Number of C-O (i.e. single bonds) bonds in carbonate ion is three.

Bondorder=TotalnumberofbondsNumberofCObonds=43=1.33\begin{array}{c}\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total}}\,{\rm{number}}\,{\rm{of}}\,{\rm{bonds}}}}{{{\rm{Number}}\,{\rm{of}}\,{\rm{C - O}}\,{\rm{bonds}}}}\\\\\,{\rm{ = }}\,\frac{4}{3}\\\\ = \,1.33\\\end{array}

Ans: Part a

Thus, the negative charge on each oxygen atom of the carbonate ion is 0.67 - {\rm{0}}{\rm{.67}}

Part b

Thus, the bond order at each carbon-oxygen bond in the carbonate ion is 1.331.33